OpenStudy (anonymous):

The equation is y=-3(x-2)^2+2. I'm supposed to graph it. I Know that the vertex is (2,2) but how do I find the x and y intercepts?

OpenStudy (anonymous):

Refer to the attachment from Mathematica.

OpenStudy (anonymous):

That's a great graph but I need to be able to do it without a graphing calculator.

hartnn (hartnn):

To get x-intercept put y=0 in the equation y=-3(x-2)^2+2 to get y-intercept put x=0 in the equation y=-3(x-2)^2+2